A QCM contains 10 questions, the student has to choose only one correct answer from 3 possible answers. How many different responses can we get?
1
Expert's answer
2020-12-24T17:14:43-0500

Let un be the number of all possible responses we can get, if QCM contains n questions.

Let's show then, that for all n un+1 = 3un.

Indeed, any response for n+1 questions consists of two parts: a response to the first n questions (total of un variants) and an answer to the last one (3 variants). All variants from the first part are compatible with any variant from the second part. Therefore, the total number of responces un+1 is a product: un * 3.

The sequence u1, u2, ..., un is a geometric progression with the common ratio 3 and, by the formula of the n-th term of it, we have un = u1*3n-1. So, u10 = 3*39 = 310 = 59049.


Answer: There exist 59049 different responses.